OpenStudy (anonymous):

Five less than the product of 7 and a number is equal to 3

OpenStudy (texaschic101):

" product " is a multiplication problem 7n - 5 = 3

OpenStudy (anonymous):

thts not what trips me up, what screws me up is when it says less than the product. can you help me wit another one?

OpenStudy (texaschic101):

sure

OpenStudy (anonymous):

twice the sum of a number and 4 is 3, would this be 2+4x=3

OpenStudy (alexandervonhumboldt2):

2*(4+n)=3 i guess

OpenStudy (texaschic101):

no... 2(4 + n) = 3

OpenStudy (anonymous):

so if it says twice or something similar it's distributing?

OpenStudy (texaschic101):

twice the sum...implies multiplying 2 times the sum (4 + n)....= 3

OpenStudy (anonymous):

ok, i realise this isn't hard, this is my kryptoninte rn for some reasone, can you help me with another one then

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (texaschic101):

ask away :)

OpenStudy (anonymous):

the difference of a number divided by 7 and 5 is equal to 2

OpenStudy (anonymous):

lemme do it

OpenStudy (alexandervonhumboldt2):

n/7 - 5=2 i tgink

OpenStudy (texaschic101):

this one is tricky

OpenStudy (anonymous):

x/(7-5)=2

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

I think so

OpenStudy (anonymous):

wait it's right?

OpenStudy (texaschic101):

I am not 100% sure...

OpenStudy (texaschic101):

its worded weird

OpenStudy (anonymous):

it's not right

OpenStudy (anonymous):

x-7/5=2?

OpenStudy (texaschic101):

hmmm....difference implies subtraction....x - 7/5 = 2...possibly

OpenStudy (texaschic101):

let me check.. @Vocaloid ...can you take a look at this please

OpenStudy (anonymous):

it's wrong

OpenStudy (alexandervonhumboldt2):

the x-7/5=2 is the differenece of a number and 7 divided by 5 equals 2

OpenStudy (alexandervonhumboldt2):

i say x/7 - 5=2

OpenStudy (anonymous):

@AlexandervonHumboldt2 i did that and it said it was wrong

OpenStudy (texaschic101):

-n / (7/5) = 2 ??.....but thats not the difference is it ?

OpenStudy (alexandervonhumboldt2):

hmmm

OpenStudy (anonymous):

no that would be the negative of a number right?

OpenStudy (texaschic101):

lol...yes...I am running out of ideas

OpenStudy (anonymous):

y is this so difficult?

OpenStudy (texaschic101):

@Directrix ...help

OpenStudy (texaschic101):

its the wording

OpenStudy (anonymous):

i realise

OpenStudy (texaschic101):

@FibonacciChick666 ..can you take a look at this

OpenStudy (texaschic101):

somebody will know this

OpenStudy (texaschic101):

n/7 - n/5 = 2 ???

OpenStudy (anonymous):

just googled it, thts the right answer

OpenStudy (texaschic101):

thats just messed up...lol...lucky guess

OpenStudy (fibonaccichick666):

ok, so what is up?

OpenStudy (anonymous):

wait no, tht was wrong, it gave me the answer, its (y/7)-5=2

OpenStudy (texaschic101):

we figured it out...the wording just messed with our brains

OpenStudy (anonymous):

i still dont get this stuff very well can we do another one?

OpenStudy (texaschic101):

oops...guess we didn't figure it out

OpenStudy (fibonaccichick666):

uh, sure but if it was the difference one the post like two up is correct

OpenStudy (anonymous):

nine more than the quotient of a number and 2 equals 3

OpenStudy (anonymous):

9(x/2)=3

OpenStudy (texaschic101):

(n/2) + 9 = 3

OpenStudy (fibonaccichick666):

close spatter, but remember more usually implies addition

OpenStudy (fibonaccichick666):

here is a really good sheet for this http://academic.venturacollege.edu/mbowen/courses/handouts/h_translate.pdf

OpenStudy (anonymous):

ok, i was thinking multiplication for some reason

OpenStudy (fibonaccichick666):

that is usually the word of

OpenStudy (texaschic101):

true..." of " means multiply..." is " means equals.....math has its own language

OpenStudy (fibonaccichick666):

and as a side note, if we are nitpicking, you have to have 9+ for the last one, not +9

OpenStudy (texaschic101):

oh yes...your right...the wording is very important in these types of problems

OpenStudy (anonymous):

i put @texaschic101 answer in and it was right so i dont think the computer is being to picky

OpenStudy (anonymous):

ok, so one more?

OpenStudy (anonymous):

the sum of a number times 4 and 5=2

OpenStudy (fibonaccichick666):

so what do you think?

OpenStudy (anonymous):

x(4+5)=2

OpenStudy (fibonaccichick666):

close, but you have the sum of( a number times four and 5) is..

OpenStudy (fibonaccichick666):

I always read these backwards to be honest

OpenStudy (anonymous):

so wait thats what it was asking for right?

OpenStudy (fibonaccichick666):

not what you wrote

OpenStudy (texaschic101):

4n + 5n = 2 ??

OpenStudy (fibonaccichick666):

what you wrote is the product of a number and the sum of four and 5 is 2

OpenStudy (fibonaccichick666):

no tex

OpenStudy (anonymous):

ahh, see i dont think of like that

OpenStudy (fibonaccichick666):

did you check out the sheet I linked to? It has some good examples of common phrases

OpenStudy (anonymous):

doin it rn

OpenStudy (texaschic101):

n(4 + 5) = 2 ??

OpenStudy (anonymous):

x+4*5=2?

OpenStudy (fibonaccichick666):

no, what you have above is the sum of a number and the product of 4 and 5 is 2

OpenStudy (fibonaccichick666):

and tex, yours is the product of a number and the sum of 4 and 5 is...

OpenStudy (anonymous):

the sum of a number = x+

OpenStudy (texaschic101):

but it says " of a number TIMES 4 and 5 "....so isn't their multiplication involved

OpenStudy (fibonaccichick666):

think of the word sum as a function. It has this format sum(______and______)=____+_____

OpenStudy (anonymous):

times 4 and 5= (4+5)=2

OpenStudy (fibonaccichick666):

there is, but not where you have it

OpenStudy (anonymous):

x+(4+5)=2

OpenStudy (fibonaccichick666):

no, that is the sum of a number and the sum of 4 and 5 is 2

OpenStudy (texaschic101):

lol....Fibonacci...your killin me here

OpenStudy (anonymous):

2+(4*5)=2?

OpenStudy (anonymous):

i'm going to drop out of HS if this isn't right lol

OpenStudy (fibonaccichick666):

"the sum of \(\color\red{\text{a number times 4}}\) and\(\color\red{ 5}\)=2"

OpenStudy (fibonaccichick666):

no, that was not correct, but don't drop out. I highlighted the necessary pieces of our function.

OpenStudy (anonymous):

(4*5)+x=2

OpenStudy (fibonaccichick666):

so the sum of ____ and ___ here is the red part. So we end up with 4n+5=2

OpenStudy (fibonaccichick666):

the and acts as a divider

OpenStudy (fibonaccichick666):

they are separate entities

OpenStudy (anonymous):

ok, im in shock tht I didn't realise that

OpenStudy (texaschic101):

my brain hurts

OpenStudy (anonymous):

six less than the quotient of a number and 5 equals 2

OpenStudy (fibonaccichick666):

I will highlight the separate entities